OpenStudy (anonymous):

List all x-intercepts for y=-3sin[(1/2)x+(1/5)pi] on the interval (-(2/5)pi,4pi)

OpenStudy (anonymous):

I'm having a hard time figuring this one out.

OpenStudy (ash2326):

We have \[y=-3 \sin {(\frac 12 x +\frac 15 \pi})\] we need to find x - intercepts so y=0 \[0=-3 \sin {(\frac 12 x +\frac 15 \pi})\] so it boils down to \[0=\sin {(\frac 12 x +\frac 15 \pi})\] do you know for what value of x sin x=0?

OpenStudy (anonymous):

At 0 and pi, right?

OpenStudy (ash2326):

these are just two values, it's 0 for n*pi n=0, 1, 2, 3,.... let's put n=0, so we get \[\frac 1 2 x+\frac 1 5 \pi=0\] from this what do you get for x?

OpenStudy (anonymous):

I'm stuck.

OpenStudy (ash2326):

\[\frac 12 x=-\frac 1 5 \pi\] \[x=-\frac 25 \pi\] does this lie in our interval?

OpenStudy (anonymous):

Yes, it does.

OpenStudy (ash2326):

we have the interval as \[(-\frac 2 5 \pi, 4\pi)\] it's open interval the end limits won't be considered

OpenStudy (anonymous):

All of the possible choices have three x-intercepts. And 4pi isn't one. How do I know how many increments to move along the x-axis before I reach 4pi?

OpenStudy (ash2326):

now put n=1 so \[\frac 1 2 x+\frac 1 5 \pi=\pi\] now find x from this we've to find x for n=2, 3 ,4 until x is greater than 4pi

OpenStudy (anonymous):

Ah, I see. So (8/5)pi should be next then, right?

OpenStudy (ash2326):

yeah, which does lie in the interval then n=2 continue doing this and do check every time that x lies in the interval

OpenStudy (anonymous):

Right, until it exceeds 4pi. I think I got it. Thanks, ash.

OpenStudy (ash2326):

you're welcome :D